OpenStudy (anonymous):

using Anti-derivative A stone is dropped off a cliff and hit the ground with a speed of 120 ft/s . a) what is the equation for acceleration a(t), velocity v(t), and displacement/ position s(t)? b) what is the height of the cliff ?

OpenStudy (amistre64):

v(t) = -32t + C = 120

OpenStudy (amistre64):

v(0) = 0; v(t) = -32t -32t = 120, when t= ?

OpenStudy (amistre64):

t = 15/4 s(t) = -16(15/4)^2 + Hi = 0, when Hi = ?

OpenStudy (amistre64):

Hi = 225 ft s(t) = -16t^2 + 225 v(t) = -32t a(t) = -32

OpenStudy (anonymous):

thank you because i had a problem with t idk i cud be a fraction so wasnt sure but thank you

OpenStudy (amistre64):

time is continuous; its doesnt have to be in discrete units